2x + 5y = 17

6x - 5y = -9


1
Expert's answer
2021-01-26T11:35:20-0500

Given:

"\begin{cases}
2x + 5y = 17\\
6x - 5y = -9
\end{cases}"


Solution:

Adding the second equation to the first one in order to cancel out y:

"2x + 6x + 5y - 5y = 17 - 9"

"8x=8"

"x=1"


Substituting x=1 into the first equation:

"2*1 + 5y = 17"

"5y = 15"

"y = 3"


Answer:

"x=1, y =3"


Check:

"\begin{cases}
2*1 + 5*3 = 17\\
6*1 - 5*3 = -9
\end{cases}"